Circular No. 7602

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001af IN MCG –04-24-1 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001af (cf. IAUC 7598), obtained by K. Rines on Mar. 23.21 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. The spectrum consists of a blue continuum with incipient P-Cyg lines of hydrogen and helium.

SUPERNOVA 2001X IN NGC 5921 A. Gal-Yam and O. Shemmer, Tel Aviv University, write: “A spectrum

(range 400–780 nm) of SN 2001X (cf. IAUC 7591, 7593) was obtained using the Wise Observatory 1-m telescope (+ FOSC spectrograph) on Mar. 22 by J. Dann. SN 2001X displays strong hydrogen Balmer lines, and the Hα line has a pronounced P-Cyg profile, similar to spectra of type-IIp supernova such as 1992H, several weeks past maximum light.”

V445 PUPPIS P. Garnavich and J. Quinn, University of Notre Dame, report that

time-resolved U-band photometry of this peculiar eruptive variable was obtained with the Vatican Advanced Technology Telescope on Mar. 16, 17, and 19 UT for two hours each night. The average time between consecutive exposures was 60 s. Contrary to the quiescence noted in January (IAUC 7565), the object showed significant variability in the ultraviolet. The star brightened at a rate of 0.16 mag/day over the observing run and showed 10-percent variations on individual nights. No periodicity is evident in the light curve. An additional set of images obtained at the Fred L. Whipple Observatory by J. Graves and P. Berlind on Mar. 26.1 shows the ultraviolet magnitude back to its Mar. 17 level of U = 10.2. Further monitoring at all wavelengths is encouraged.

Visual magnitude estimates: Feb. 20.403 UT, 9.8 (J. Bedient, Honolulu, HI); 27.559, 10.2 (A. Pearce, Nedlands, W. Australia); Mar. 6.783, 10.3 (M. Lehký, Hradec Králové, Czech Republic); 14.361, 9.6 (M. Linnolt, Honolulu, HI); 19.271, 9.9 (Bedient).

1998 SF_{36}

Corrigendum. On IAUC 7598, item by Hicks et al., line 6, for (V − X = +0.19 ± 0.003) read (V − X = +0.019 ± 0.003)

2001 March 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green